# Automated glenoid bone loss measurement and segmentation in CT scans for pre-operative planning in shoulder instability

### v1

@@ 2.6 Experimental Plan
To evaluate the performance of the proposed deep learning pipeline for glenoid bone measurement, we designed a set of experiments focusing on (i) segmentation accuracy, (ii) accuracy of the en-face view plane computation, (iii) accuracy of curve on the intact posterior inferior contour of the glenoid on en-face view, and (iv) accuracy of the measurement of glenoid bone loss. We tested the result predicted by the pipeline on our test data set across three different levels of bone loss: $13.5\%$, $13.5\%<\text{loss}<20\%$, and $20\%$.
Experiment 1: Glenoid Articular Surface and Humerus Bone Segmentation
Segmentation of the articular surface of the glenoid and humerus bone is compared between the prediction of the segmentation model and manual measurements of two clinicians in three different levels of bone loss. The dice coefficient is used as the accuracy metric.
Experiment 2: Curve on the Intact Contour of the Glenoid
This experiment compares the chamfer scores of the point sets of glenoid rim generated by the algorithm and the mean point sets of two clinicians.
Experiment 3: En-face View
This experiment compares the angular difference between the algorithm-predicted en-face view and consensus direction to the half angular difference between two doctors. To make a fair comparison, we first compute angular differenece between two doctors $\phi=\arccos(\hat{d}_{1}\cdot\hat{d}_{2})$, then the human error would be $\approx\phi/2$ because a single doctor’s error is $\phi/2$ from $\hat{\mu}$. Algorithm error involves the angular difference between the direction vector predicted by the algorithm and a consensus unit direction from the two doctors, same direction vector by $\phi=\arccos(\hat{d}_{1}\cdot\hat{d}_{2})$, but in a different form, $\hat{\mu}=\frac{\hat{d}_{1}+\hat{d}_{2}}{\|\hat{d}_{1}+\hat{d}_{2}\|}\quad$. Then the algorithm’s angular error would be $\theta_{A}=\arccos\!\left(\operatorname{clip}(\hat{a}\cdot\hat{\mu},-1,1)\right).$
Experiment 4: Glenoid Bone Loss
This experiment evaluated algorithm-predicted glenoid bone loss against physician measurements, using human-to-human measurements as the baseline, throughout the whole patient cohort, and within three subgroups of severity of bone loss. We used mean absolute error (MAE) to quantify error magnitude, the Pearson correlation coefficient for correlation analysis, a two-way mixed-effects, absolute-agreement, single-measurement intraclass correlation coefficient (ICC) for agreement, Bland–Altman analysis to assess systematic bias, and a confusion matrix to evaluate the model’s classification performance.
@@ 3 RESULTS
We conducted four experiments that evaluated the automated glenoid bone loss pipeline in our test data set.
@@ 3.2 Experiment 2: Evaluation of Predicted Rim on the Intact Posterior Contour of the Glenoid
We evaluated the accuracy of the rim points predicted by Rim-UNet using the bidirectional Chamfer distance between the predicted and ground-truth glenoid rim skeletons (in mm). The Chamfer distance estimates how similar the predicted rim point set is to the set annotated by the clinician on the glenoid surface. The mean nearest-neighbor distance was $1.37\pm 0.58\,\text{mm}$, with most cases having distances $1.17\,\text{mm}$.
@@ 3.3 Experiment 3: Assessment of En-face Direction Vector
We evaluated the en-face view of the glenoid computed by the algorithm against manual selections from two clinicians. The mean of a doctor’s angular error to the consensus on en face view is $4.94^{\circ}\pm 1.86^{\circ}$ (95% CI: $4.09^{\circ}$$5.78^{\circ}$), while the algorithm’s angular error to consensus is $7.85^{\circ}\pm 3.99^{\circ}$ (95% CI: $6.03^{\circ}$$9.67^{\circ}$). While the mean error of algorithm-consensus pair is higher, of 21 observation analyzed, 24% demonstrated algorithm performance below parity ($y\leq x$), and 71% fell within the clinically acceptable non-inferiority margin of $5^{\circ}$ ($y\leq x+\Delta\theta$).
@@ 4 DISCUSSION
The most important finding in this study is that (1) a computer vision and deep learning-based algorithm pipeline can be used to reliabily automate the measurement of glenoid bone loss on 3D CT scans. (2) Our algorithm and pipeline can be a clinically reliable tool in assisting pre-operative planning in shoulder instability, especially in screening high bone loss (¿20%) as a classification model. (3). The segmentation model in the pipeline is also clinically reliable in subtracting the humerus bone and preparing the glenoid bone segmentation for glenoid bone-related assessment. Additionally, the current automated pipeline can reliably identify true en-face view.
@@
In our analysis of bone loss, our computer vision and deep learning-based pipeline demonstrated higher agreement and correlation with fellowship-trained orthopaedic surgeon measurements compared to the inter-rater baseline between surgeons. It is much better than the human baseline in the high bone loss group (ICC 0.83 vs. 0.22) in reliability. It shows minimal systematic bias and a smaller spread of differences than the human baseline, achieving 71.4 and 85.7 in classifying low and high bone loss groups, respectively, and without any instances of misclassifying low as high or vice versa. The correlation and classification in the moderate bone loss group are weak, but are consistent with the human baseline. All of which suggests that our computer vision-based deep learning method can reliably automate the measurement of glenoid bone loss in CT scans. Our pipeline is clinically reliable for quantitative measurements and effectively identifies cases of high bone loss, although we recommend human review for cases within the 13.5% to 20% ”gray zone.” It is important to note that in the gray zone- that clinical surgeon decision making remains gray as well. The current algorithm can trigger surgeons that bone loss is within this zone and requires greater scrutiny as would be the case with manual measurements. In areas of large bone loss- the algorithm accurate measures bone loss which can allow surgeons to reliability plan bony augmentation techniques.
@@
The humerus bone segmentation dice score of $0.9866\pm 0.0059$ demonstrated that our segmentation model can be an excellent tool for humerus bone subtraction and, therefore, a great tool to help visualize the glenoid bone. As shown in Figure 1, the humerus bone is completely removed, leaving no residual bone fragment and a complete view of the glenoid bone. It is important to note- that current clinical CT 3D protocols require manual segmentation to remove humerus and create 3D rendered glenoid/scapula images. Based on the reliable bone loss result demonstrated in the pipeline in this study and the fact that the pipeline is clinically reliable in analyzing glenoid bone loss, we believe it also proves that the segmentation model, which is an essential part of the pipeline, is also reliable in helping to make clinical decisions. This segmentation model can be a valid tool for pre-operative planning in shoulder instability.
A lower glenoid bone segmentation score does not necessarily mean that the segmentation model performance is poorer in the glenoid bone. However, it could mean that we do not have a better segmentation metric for evaluating just part of the glenoid bone on the surface toward the joint side that we emphasize. There is a 5% favor in the humerus bone dice score. This difference is expected, as the humerus bone has an enclosed structure that can be easily annotated. In contrast, the glenoid articular surface has a much gray zone defining the opposite side to the articular glenoid surface, where doctors will not have a clear consensus. However, this ambiguity on the other side does not affect the accuracy of the pipeline, as we would only be using the actual side for rim points set finding (see Fig. 1
The chamfer distance of the rim points between the algorithm’s prediction and the doctor’s consensus indicates strong agreement between the predictions and the clinician’s labels for the rim location on the glenoid surface. Most patients show a distance of less than 1.17 mm, which is minimal compared to the surface area of the glenoid. This shows a good foundation for fitting the circle to these points on the rim for the prediction of bone loss.
Finding an en-face view is challenging, as there is no clear definition of it in clinical settings. We defined the mean of two fellowship trained surgeons directions as consensus and compared the angular error of both the single doctor and the algorithm. Although only 24% of patients show lower error than a single doctor, 71% fell within the clinically acceptable non-inferiority margin of $5^{\circ}$ ($y\leq x+\Delta\theta$), suggesting that while the algorithm generally exceeds inter-observer variability, it remains within a clinically tolerable range for the majority of cases.
4.1 Why not PCA / largest projected area?
4.2 Rationale for our approach
One thing to address is Rim-Unet’s ability to learn the representation of the en-face view and the glenoid rim for circle fitting within a single model. We made this possible because of our data collection method. We first ask the surgeon to reorient the glenoid articular surface to the en-face view, which is normal to the screen. Then, on this plane, surgeons’ are required to select the rim points containing the plane’s spatial information. Therefore, the models are trained to find rim points already containing the en-face view information. To illustrate this concept, imagine drawing the contour of a bottle in a 3D space from a fixed angle. The paint on that sphere is actually drawn on a 2D space because we are forced to draw on the edge of the object from one angle; therefore, we can utilize the points on that paint to find the en-face direction because the points on the contour should be orthogonal to that angle we were facing.

### v2

@@ 2.5 Evaluation
To evaluate the proposed deep-learning pipeline for glenoid bone measurement, we designed experiments focusing on (i) segmentation accuracy, (ii) accuracy of en-face view plane computation, (iii) accuracy of the fitted curve along the intact posterior-inferior glenoid contour on the en-face view, and (iv) accuracy of the final glenoid bone-loss measurement. We evaluated pipeline predictions on the test set across three bone-loss severity levels: $13.5\%$, $13.5\%<\text{loss}<20\%$, and $20\%$.
@@ 3.2 Experiment 2: Evaluation of Predicted Rim on the Intact Posterior Contour of the Glenoid
We evaluated the accuracy of the rim points predicted by RimU-Net using the bidirectional Chamfer distance between the predicted and ground-truth glenoid rim skeletons (in mm). The Chamfer distance estimates how similar the predicted rim point set is to the set annotated by the clinician on the glenoid surface. The mean nearest-neighbor distance was $1.37\pm 0.58\,\text{mm}$, with most cases having distances $1.17\,\text{mm}$.
@@ 3.3 Experiment 3: Assessment of En-face Direction Vector
We evaluated the en-face view of the glenoid computed by the algorithm against manual selections from two clinicians. The mean of a doctor’s angular error to the consensus on en face view is $4.94^{\circ}\pm 1.86^{\circ}$ (95% CI: $4.09^{\circ}$$5.78^{\circ}$), while the algorithm’s angular error to consensus is $7.85^{\circ}\pm 3.99^{\circ}$ (95% CI: $6.03^{\circ}$$9.67^{\circ}$). While the mean error of algorithm-consensus pair is higher, of 21 observation analyzed, 24% demonstrated algorithm performance below parity ($y\leq x$), and 71% fell within the clinically acceptable non-inferiority margin of $5^{\circ}$ ($y\leq x+\Delta\theta$)(Fig 2).
@@ 4 Discussion
The most important finding in this study is that (1) a fully automated deep learning-based algorithm pipeline can be used to reliabily automate the measurement of glenoid bone loss on 3D CT scans. (2) Our algorithm and pipeline can be a clinically reliable tool in assisting pre-operative planning in shoulder instability, especially in screening high bone loss ($20\%$) as a classification model. (3). The segmentation model in the pipeline is also clinically reliable in subtracting the humerus bone and preparing the glenoid bone segmentation for glenoid bone-related assessment. Additionally, the current automated pipeline can reliably identify true en-face view.
@@
In our analysis of bone loss, our automated pipeline demonstrated higher agreement and correlation with fellowship-trained orthopaedic surgeon measurements compared to the inter-rater baseline between surgeons. It is much better than the human baseline in the high bone loss group (ICC 0.83 vs. 0.22) in reliability. It shows minimal systematic bias and a smaller spread of differences than the human baseline, achieving 71.4 and 85.7 in classifying low and high bone loss groups, respectively, and without any instances of misclassifying low as high or vice versa. The correlation and classification in the moderate bone loss group are weak, but are consistent with the human baseline. All of which suggests that our deep learning-based algorithm can reliably automate the measurement of glenoid bone loss in CT scans. Our pipeline is clinically reliable for quantitative measurements and effectively identifies cases of high bone loss, although we recommend human review for cases within the 13.5% to 20% ”gray zone.” It is important to note that, in the gray zone, clinical surgeon decision-making remains gray as well. The current algorithm can alert surgeons that bone loss is within this zone and requires greater scrutiny, as would be the case with manual measurements. In areas of large bone loss, the algorithm accurately measures bone loss, which can allow surgeons to reliably plan bony augmentation techniques.
@@
The humerus bone segmentation dice score of $0.9866\pm 0.0059$ demonstrated that our segmentation model can be an excellent tool for humerus bone subtraction and, therefore, a great tool to help visualize the glenoid bone. As shown in Figure 1, the humerus bone is completely removed, leaving no residual bone fragment and a complete view of the glenoid bone. It is important to note that current clinical CT 3D protocols require manual segmentation to remove humerus and create 3D rendered glenoid/scapula images. Based on the reliable bone loss measurements demonstrated by this pipeline and its overall clinical reliability for assessing glenoid bone loss, we believe the segmentation model, which is an essential component of the pipeline, is also reliable for supporting clinical decision making by providing glenoid visualization. This segmentation model may serve as a useful tool for preoperative planning in patients with shoulder instability.
@@ 5 CONCLUSION
Appendix A Appendix A
Figure 6: Flowchart showing the datasets used to train and test the automated pipeline for measuring glenoid bone loss.
A.1 Ground-Truth Generation
[…]
A.2 Two stages of fine-tuning
In the second stage, we further fine-tuned the model on our sub-dataset to capture finer substructures, as described on the glenoid articular surface[Chlebus2018]. We collected a smaller dataset under the guidance of an orthopedic surgeon, comprising 98 CT scans and their corresponding left/ right glenoid articular surfaces and humerus bone loss, as shown in Table 1. The training plan used to fine-tune the first-stage model is identical to the one used in the first stage. We used the same loss function for both stages, L = LCE + LDICE, which combines entropy loss and the dice loss. We will release the exact plan/config JSON, trained weights, and scripts for full reproducibility.
Appendix B Appendix B
The RimU-Net is adapted from the structure of a vanilla U-net tailored for the heat-map regression thin-curve task. In addition to a basic encoder-decoder structure, we used strided convolution $(2,2,1)$ to replace the max-pooling for downsampling in the first two stages to preserve the details of the plane for our anistropic medical images [Isensee2020]. Moreover, for each convolution, we use Instance Normalization for medical images in small batches [Verweij2020] and LeakyReLU activations with $\alpha=0.01$ to preserve all gradients flowing to the prediction of our sparse target output [Maas2013]. We used adaptive wing loss for our landmark localization task, as a greater penalty is needed for medium to small errors. [Feng2018Wing][Wang2019AdaptiveWing].
[…]
Appendix C Appendix C
C.1 Why not PCA / largest projected area?
C.2 Rationale for our approach
One thing to address is RimU-Net’s ability to learn the representation of the en-face view and the glenoid rim for circle fitting within a single model. We made this possible because of our data collection method. We first ask the surgeon to reorient the glenoid articular surface to the en-face view, which is normal to the screen. Then, on this plane, surgeons’ are required to select the rim points containing the plane’s spatial information. Therefore, the models are trained to find rim points already containing the en-face view information. To illustrate this concept, imagine drawing the contour of a bottle in a 3D space from a fixed angle. The paint on that sphere is actually drawn on a 2D space because we are forced to draw on the edge of the object from one angle; therefore, we can utilize the points on that paint to find the en-face direction because the points on the contour should be orthogonal to that angle we were facing.